OpenStudy (anonymous):

Write an equation of the line that passes through (2,-7) and is parallel to the line x=5.

OpenStudy (anonymous):

x=2 :)

OpenStudy (anonymous):

Can you please show me how it is done? Thanks! :)

OpenStudy (anonymous):

because if its parallel to x=5 then its vertical, which means it will be x=?, and in this case, x=2!!

OpenStudy (anonymous):

I need to show work, I need to see steps to actually solving the problem for me to actually understand it well.

OpenStudy (anonymous):

 this shows the graph of x=5

OpenStudy (anonymous):

if its parallel to that, then it must also be veritcal, since parallel lines never touch.

OpenStudy (shoba):

use this equation y-y1 = m(x-x1) (x1,y1) = (2,-7) x =5 ------x/5 = 0 m=1/5 y = mx + c then substitude into that equation at point (2,-7) y+7 = 1/5(x-2) y = x/5 - 2/5 - 7 y= x/5 -33/5

OpenStudy (anonymous):

 the second line is vertical AND passes through 2,-7, and the equation for it is x=-2. since its vertical the equation is x=?, and in this case, x=-2!

OpenStudy (anonymous):

@Shoba, can you please show me how doing that gives me x=-2?